May 4, 2016 Calibration

From GlueXWiki
Jump to: navigation, search

GlueX Calibration Meeting
Wednesday, May 4, 2016
11:00 am, EST
JLab: CEBAF Center, F326

Communication Information

Remote Connection

You can connect using BlueJeans using the meeting number 630 804 895 .       (Click "Expand" to the right for more details -->):

  1. Make sure you have created a BlueJeans account via your JLab CUE account using this link:

  2. Meeting ID: 630804895
    • (you may need to type this in, depending how you connect)

  3. If connecting via Web Browser: click this link (no passcode is needed):

  4. If connecting via iOS or Android App:
    • Use your JLab e-mail address to log in and then enter the meeting ID given above to join the meeting

  5. If connecting via Phone: Dial one of the following numbers and then enter the meeting ID above and hit "#" or "##"

  6. If connecting via Polycom unit:
    • Dial or
    • Enter meeting ID above
    • Use *4 to unmute


Talks can be deposited in the directory /group/halld/www/halldweb/html/talks/2016 on the JLab CUE. This directory is accessible from the web at .


  1. Announcements
  2. Collaboration Meeting
  3. Data Production Planning
  4. Subdetector Reports
  5. Calibration Tasks
  6. Simulations
  7. AOB


Attending: Sean (NU); Simon, Adesh, Eugene, Mark I., Alex B., Luke; Curtis, Mike S., Will M., Naomi (CMU); Mahmoud (FIU)

  1. Announcements
    • Final calibration run is underway, delayed for ~3 days by volatile disk problems, skim jobs are hitting the farm, ~10 % done, should be finished by end of next week. Sean is fixing some calibrations, expected to be done tomorrow.
    • New GitHub repo was created for non-sim-recon calibration software hd_calibration.
  2. Collaboration Meeting is next week!
  3. Sean went over the to-do items for the first reconstruction pass planned at the end of the month. Eugene made several comments that encouraged some discussion.
    • Preliminary relative PS fluxes based on coincidences between the two arms are planned to be added to the CCDB. Eugene pointed out that the PS data needs to be analyzed in more detail, as a function of time during the run, etc. The PS triggers are being skimmed, but it's unclear who is going to do this job. We may only want to keep physics events during certain parts of the run. The PS reconstruction in sim-recon is very bare-bones and needs to be improved, and improvements to the CCDB would be needed to keep information that changes during the run.
    • It would be useful to keep statistics of the different triggers for all of the data. Sean pointed out that Elton has some code that could be used for this purpose, and will work with him to adapt this for standard monitoring purposes. One still needs to decode the meaning of these triggers.
    • We need to be sure that we're only keeping physics triggers in standard analysis code.
  4. Subdetector Reports
    • FCAL - Adesh reported that he is still studying the gain calibrations and energy non-linearities. Jon Z. is working on efficiency studies.
    • BCAL - Will is finishing up fixing some 2nd order effects in the new clustering code and the shower gain/energy calibration from the last run. Mark D., Will, and George are making progress on refining the BCAL timing. Mark I. offered to help Will set up a project account for his calibration jobs. The farm is busy with calibration and monitoring jobs, so the project accounts are practically the only ones getting processing time.
    • Tracking - Simon has pushed his geometry changes to a branch of HDDS and doing some final checks.
    • TOF - Beni has been updating the CCDB with new timing and gain calibrations.
    • Start Counter - Sean showed some plots that he'd shown at the last Start Counter meeting, which better shows the mismatch between the projected track and nominal start counter hit location, which has been seen in other analyses. Simon is working on defining a looser matching cut in the nose region, better alignment is the longer term solution. Sean has started looking at similar plots for the TOF.
    • TAGM - Alex B. has been iterating between improving the timing calibrations on the microscope and determining software thresholds for analysis. He is mostly done with the data after the TAGM gains were changed towards the end of the run, and is applying similar techniques to improve the earlier data. Using PS triggers, he is getting a preliminary efficiency of ~75%, in line with expectations. He noted that the running with the 2ns bunch spacing was a problem for the poor quality fibers, since they have poorer resolution. This shouldn't be a problem with the new fibers being manufactured.
    • CDC - Naomi took a first look at the CDC gain calibrations, and the charge distribution as a function of wire showed some odd periodic-ish structures. These could be correlated with the location of the straws. In any case, this looks like it won't be a trivial project.